OpenStudy (anonymous):

1. Evaluate the following expression for m = 36. √ 4m + 5

OpenStudy (anonymous):

\[\sqrt{4(36)+5}\] \[\sqrt{144+5}\] \[\sqrt{149}\] \[=12.2066\]